y varies jointly as x and z and inversely as the product of w and p, and y=45/160 when x=5,z=9,w=5, and p =8

the directions are to find an equation of variation in which the following are true.

Answer by stanbon(75887)

y varies jointly as x and z and inversely as the product of w and p
---
y = k[xz/wp]
----------------------
and y=45/160 when x=5,z=9,w=5, and p =8
---
(45/160) = k[5*9/5*8]
(45/160) = k[9/8]
k = (8/9)(45/160)
k = 5/20
k = 1/4
-------------
Therefore the equation is
y = (1/4)[xz/wp]
=========================
Cheers,
Stan H.